OpenStudy (anonymous):

If f(x) = 3x − 2 and g(x) = 2x − 1, find (f − g)(2). 7 1 −7 −1

OpenStudy (anonymous):

find \(f(2)\) then find \(g(2)\) then subtract

OpenStudy (anonymous):

do you know how to find \(f(2)\) ?

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

Perhaps easier would be to notice that \[\eqalign{f(x)-g(x)&=(3x-2)-(2x-1) \\ &=3x-2x-2+1 \\ &=x-1 \\}\] So \((f-g)(2)=(x-1)\bigg|_{x=2}=(2)-1=2-1=1\)

OpenStudy (zzr0ck3r):

that does not seem easier to me:) f(2) = 4 g(2) = 3 4-3=1